OpenStudy (anonymous):

y = –x + 2 y = 3x – 1 =( help please

OpenStudy (anonymous):

Is y = 5/4?

OpenStudy (anonymous):

ummm not sure haha im just learned this like 15 mins ago,

OpenStudy (anonymous):

Solving Systems by Graphing

OpenStudy (anonymous):

Well I m not sure, but I think x = 3/4 & y= 5/4

OpenStudy (anonymous):

umm okay ill try it.